# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v2

@@ Abstract
Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention. However, existing link prediction models are usually designed for one specific type of facts, making it difficult to generalize to other fact representations. To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction. It consists of a unified Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module as graph encoder. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts. Experimental results across 7 datasets from 3 types of KGs demonstrate that our UniHR outperforms baselines designed for one specific kind of KG, indicating strong generalization capability of HiDR form and the effectiveness of HiSL module. Code and data are available at https://github.com/Lza12a/UniHR.
UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction
Zhiqiang Liu1,3, Mingyang Chen2,3, Yin Hua1,3, Zhuo Chen2,3, Ziqi Liu4, Lei Liang4, Huajun Chen2,3, Wen Zhang1,3thanks: Corresponding Author. 1School of Software Technology, Zhejiang University 2College of Computer Science and Technology, Zhejiang University 3ZJU-Ant Group Joint Lab of Knowledge Graph 4Ant Group {zhiqiangliu,zhang.wen}@zju.edu.cn
@@ 1 Introduction
Large-scale knowledge graphs (KGs) such as WordNet Miller (1995), Freebase Bollacker et al. (2008), and Wikidata Vrandečić and Krötzsch (2014) have been widely applied in many areas like question answering Kaiser et al. (2021), recommendation systems Guo et al. (2020), and natural language processing Annervaz et al. (2018). However, the presence of missing facts within these KGs inevitably limit their applications. Therefore, the link prediction task has been introduced to predict missing elements within factual data. Current link prediction methods mainly focus on facts in the form of triple $\left(head\,entity,relation,tail\,entity\right)$.
Figure 1: A special KG consists of triple-based fact, hyper-relational fact, nested fact and temporal fact.
@@
Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1. Therefore, existing researches Wang et al. (2021); Xiong et al. (2024); Xu et al. (2019) contribute to focusing on semantically richer facts. Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation. Thus in recent years, Hyper-relational KGs (HKG) Chung et al. (2023), Temporal KGs (TKG) Xu et al. (2023a), and Nested factual KGs (NKG) Xiong et al. (2024) attract wide research interests.
@@
Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations Xiong et al. (2023). However, these methods are usually designed for specific representation forms, e.g., StarE Galkin et al. (2020) customizes graph neural network to implement message passing on hyper-relational facts, For nested factual KGs, BiVE Chung and Whang (2023) connects two levels of facts throgh a simple linear layer. In addition, GeomE+ Xu et al. (2023a) et al. temporal KG embedding methods contain time-aware scoring functions to adapt timestamps. Although these methods perform well on specific type of facts, it is evident that such customized methods are difficult to generalize to other types of KGs. Therefore, establishing a unified representation learning method for multiple types of KGs is worth to investigate.
@@
To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global view of nodes to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Experimental results demonstrate that our UniHR achieves state-of-the-art performance on HKG and NKG datasets, and competitive performance on TKG datasets, revealing strong generalization capability of HiDR form and effectiveness of HiSL module. Our contributions can be summarized as follows:
@@
    We propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
@@
    We conduct link prediction experiments on 7 datasets across 3 types of KGs. Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
@@ 2 Preliminaries
Triple-based Knowledge Graph.
A common triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations.
@@
Link Prediction on Triple-based KG. The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
Hyper-relational Knowledge Graph.
A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$. Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs.
@@
Link Prediction on Hyper-relational KG. Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs. Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ ((?$,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
Nested Factual Knowledge Graph.
A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts. =$\mathcal{F}\,\text{=}\left\{\left(h,r,t\right)|h,t\in\mathcal{V},r\in\mathcal{R}\right\}$ is the set of atomic facts, where $\mathcal{V}$ is a set of atomic entities and $\mathcal{R}$ is a set of atomic relations. ^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations.
@@
Link Prediction on Nested Factual KG. The link prediction on the NKGs is performed on the atomic facts or nested facts. We refer to the link prediction on atomic facts as Base Link Prediction, and the link prediction on nested facts as Triple Prediction. For base link prediction, given a query $\left(h,r,?\right)$ or (?$,r,t\right)$, the aim is to predict the missing atomic entity ‘????’ from $\mathcal{V}$. For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
Temporal Knowledge Graph.
@@
A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps.
@@
Link Prediction on Temporal KG. The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
@@ 3 Related Works
Earlier HKG representation learning methods e.g., m-TransH Wen et al. (2016), RAE Zhang et al. (2018) have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs. Galkin et al. customize StarE Galkin et al. (2020) based on CompGCN Vashishth et al. (2019) for hyper-relational facts to capture the interaction information of key-value pairs with the main triple in the message passing stage, and achieves impressive results, demonstrating that the structural information of the graph in HKGs is also important. GRAN Guan et al. (2021) introduces edge-aware bias into the vanilla transformer attention Vaswani et al. (2017), while HyNT Chung et al. (2023) designs a qualifier encoder for HKG. They both focus on intra-fact dependencies but ignore the global structural information. Due to the existence of its particular key-value pairs on H-Facts, there are many limitations in capturing global structural information through the widely available triple-based GNNs.
@@
Chung et al. Chung and Whang (2023) first introduced the concept of nested facts. They also propose BiVE which projects atomic facts to fact nodes in the encoding phase via a simple linear layer and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE Zhang et al. (2019) or BiQUE Guo and Kok (2021). Based on BiVE, NestE Xiong et al. (2024) represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions. These methods only capture the semantic information between atomic facts and nested facts while ignoring global structural information. Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact.
@@
Recent studies in temporal knowledge graph representation learning have focused on enhancing performance by designing special time-aware scoring functions. Models such as TTransE Leblay and Chekol (2018), HyTE Dasgupta et al. (2018), TeRo Xu et al. (2020), and TGeomE+ Xu et al. (2023a) incorporate temporal-aware module into the KGE score function in various ways. However, existing models for link prediction seldom directly utilize GNNs to perceive time information for enhancing entity and relation embeddings. Therefore, we believe it is a direction worth exploring.
@@ 4 Methodology
In this section, we introduce our method, a Unified Hierarchical Representation learning framework (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form. 2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module with the enhancement of semantic information within individual facts and structural information between facts on the whole graph. 3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
Figure 2: Diverse facts are translated into the HiDR form. Blue, light blue, and red circles denote atomic nodes, relation nodes, and fact nodes. Black, light blue, and red arrows denote atomic relations, connected relations, and nested relations. The triples they connect correspond to atomic facts, connected facts, and nested facts.
@@ 4.1 Hierarchical Data Representation
To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR. Different from labelled RDF representation Ali et al. (2022), we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
As shown in Fig. 2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR. Inspired by the nested fact form Xiong et al. (2024), we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node. To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$. These relation nodes are derived from transforming the relation edges in the original KG. It is important to facilitate direct access of fact nodes to the relevant atomic nodes during message passing process. To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes. Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and a atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
Definition 1.
Hierarchical Data Representation:
A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$). $\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,$ $has\,tail\,entity\}$. The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.

### v3

@@ Abstract
Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention. However, constrained by complex fact representation forms, existing link prediction models for beyond-triple facts have difficulty achieving hierarchical fact modeling and generalizing the modules for one specific facts to other fact types. To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction. It consists of a unified Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module as graph encoder. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts. Empirical results demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations. Code and data are available at https://github.com/Lza12a/UniHR.
Keywords:
Knowledge graph representation learning Beyond-triple knowledge graph Link prediction.
@@ 1 Introduction
Large-scale knowledge graphs (KGs) such as Freebase [3] and Wikidata [33] have been widely applied in many areas like question answering [17] and natural language processing [2]. However, the presence of missing facts within these KGs inevitably limit their applications. Therefore, the link prediction task has been introduced to predict missing elements within factual data. Current link prediction methods mainly focus on the fact in the form of triple e.g., $(head\,entity,relation,$
$tail\,entity)$.
Figure 1: Comparison between the existing link prediction methods for specific beyond-triple knowledge graphs and our unified methods.
@@
Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1. Therefore, existing researches [34, 37, 38] contribute to focusing on semantically richer facts. Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation. Thus in recent years, Hyper-relational KGs (HKG) [6], Temporal KGs (TKG) [40], and Nested factual KGs (NKG) [37] attract wide research interests.
@@
Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations [36]. However, these methods are hard to achieve hierarchical fact modeling and usually designed for specific representation forms, e.g., StarE [12] customizes graph neural network to enhance inter-fact structure information between hyper-relational facts. For NKGs, BiVE [7] bridges intra-fact semantics between two levels of facts through a simple MLP. Moreover, GeomE+ [40] et al. TKG methods employ time-aware scoring functions to capture intra-fact collaborative information between factual semantics and temporal information in the timestamp. Although these methods perform well on specific types of facts, it is evident that such customized methods are difficult to generalize to other types of KGs for intra-fact structure and inter-fact semantic enhancements. Therefore, establishing a unified hierarchical representation learning method for multiple types of KGs is worth to investigate.
@@
To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Empirical results demonstrate the UniHR achieves state-of-the-art or competitive performance on most metrics and highlight the strong potential of unified representations. Our contributions can be summarized as follows.
@@
    To our knowledge, we propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
@@
    We conduct link prediction experiments on 9 datasets across 5 types of KGs. Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
@@ 2 Preliminaries
Link Prediction on Triple-based KG.
@@
A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations. The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
Link Prediction on Hyper-relational KG.
@@
A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$. Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs. Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs. Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ ((?$,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
Link Prediction on Nested Factual KG.
@@
A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts. =$\mathcal{F}\,\text{=}\left\{\left(h,r,t\right)|h,t\in\mathcal{V},r\in\mathcal{R}\right\}$ is the set of atomic facts, where $\mathcal{V}$ is a set of atomic entities and $\mathcal{R}$ is a set of atomic relations. ^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations. The link prediction on the NKGs is performed on the atomic facts or nested facts. We refer to the link prediction on atomic facts as Base Link Prediction, and the link prediction on nested facts as Triple Prediction. For base link prediction, given a query $\left(h,r,?\right)$ or (?$,r,t\right)$, the aim is to predict the missing atomic entity ‘????’ from $\mathcal{V}$. For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
Link Prediction on Temporal KG.
@@
A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,$
@@
$[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps. The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
@@ 3 Related Works
Earlier HKG representation learning methods e.g., m-TransH [35], RAE [44] have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs. Galkin et al. first customize StarE [12] based on CompGCN [31] for H-Facts to capture global structure information between H-Facts in the message passing stage, and achieves impressive results, demonstrating that the structure information of the graph in HKGs is also important. GRAN [15] introduces edge-aware bias into the vanilla transformer attention [32], while HyNT [6] designs a qualifier encoder for HKG. They both focus on intra-fact semantic dependencies but ignore the global structure information. In order to fully capture inter-fact and intra-fact information, HAHE [23] first adopts dual graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling and achieves significant performance improvement. We consider the hierarchical attention mechanism of HAHE a good beginning, but it is customized for HKG and is difficult to generalize to other facts.
@@
Chung et al. [7] first introduced nested facts to imply relationships between facts. They also propose BiVE which bridges semantics between atomic facts and fact nodes in the encoding phase via a simple MLP and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE [45] or BiQUE [16]. Based on BiVE, NestE [37] represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions. These methods only capture the intra-fact semantic information between atomic facts and nested facts. Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact for capturing inter-fact global information.
@@
Recent studies in temporal knowledge graph representation learning have focused on designing special time-aware scoring functions for inter-fact semantic modeling. Models such as TTransE [21], HyTE [8], TeRo [39], and TGeomE+ [40] incorporate temporal-aware module into the KGE score function in various ways. However, existing models seldom regard timestamps as node and directly utilize inter-fact message passing mechanism to perceive time information for enhancing entity and relation embeddings. Therefore, we believe it is a direction worth exploring.
@@ 4 Methodology
In this section, we introduce our method, a Unified Hierarchical Representation learning framework (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form. 2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph. 3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
Figure 2: Diverse facts are translated into HiDR form.
@@ 4.1 Hierarchical Data Representation
To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR. Different from labelled RDF representation [1], we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
As shown in Fig. 2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR. Firstly, we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node. To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$. These relation nodes are derived from transforming the relation edges in the original KG. It is important to facilitate direct access of fact nodes to the relevant atomic nodes during message passing process. To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes. Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
Definition 1
Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,$
$\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$). $\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,$
$tail\,entity\}$. The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.